### 18 Mar 2024 v1

# Conformally symplectic structures and the Lefschetz condition

Abstract.
This short note provides a symplectic analogue of Vaisman’s theorem in complex geometry. Namely, for any compact symplectic manifold satisfying the hard Lefschetz condition in degree 1, every locally conformally symplectic structure is in fact globally conformally symplectic, whenever there is a mutually compatible almost complex structure.
The first author is supported by the Simons Foundation Grant #636075. The second author acknowledges support provided by the Simons Foundation Award program for Mathematicians (#963783) and a PSC-CUNY Award, jointly funded by The Professional Staff Congress and The City University of New York (TRADB # 66358-00 54)
This short note provides a symplectic analogue of Vaisman’s foundational result for complex manifolds, which states that if a complex manifold is Kähler, then every locally conformally Kähler structure is indeed globally conformally Kähler [20].
Recall that a locally conformally symplectic structure is a pair $(\eta,\theta)$ consisting of a real closed $1$-form $\theta$, and a non-degenerate $2$-form $\eta$, such that $d\eta=\theta\wedge\eta$, where $d$ is the exterior derivative. If $\theta$ is $d$-exact, then $(\eta,\theta)$ is globally conformally symplectic. Moreover, if $\theta=0$, then $\eta$ is a symplectic form. For more about locally conformally symplectic structures, we refer the reader for instance to [10, 17, 2]. An almost complex structure $J$ is compatible with $\eta$ if $\eta(J\cdot,J\cdot)=\eta$ and $\eta(\cdot,J\cdot)$ defines a Riemannian metric. The induced Riemannian metric is called an almost-Kähler metric if $\eta$ is a symplectic form.
Symplectic manifolds certainly can have locally conformally symplectic structures which are not globally conformally symplectic, c.f Example 3 below (we also refer the reader to [3, 4, 1]). A sufficient condition for the desired implication, in the presence of a compatible almost complex structure, is the hard Lefschetz condition in degree $1$.
Theorem 1.
If a compact symplectic manifold $(M,\omega)$ satisfies hard Lefschetz in degree $1$, then any locally conformally symplectic structure $(\eta,\theta)$ on $(M,\omega)$ is necessarily globally conformally symplectic whenever $\eta$ and $\omega$ share a compatible $J$.
Recall that on a symplectic manifold $(M,\omega)$ of dimension $2n$ the hard Lefschetz condition holds in degree $1$ if
$L^{n-1}:H^{1}_{d}(M)\to H^{2n-1}_{d}(M)$
is an isomorphism, where $L\alpha=\omega\wedge\alpha$. We begin with a lemma, after establishing some notation.
For an almost complex manifold, we let $d_{c}:=J^{-1}dJ$, where $J$ is extended as an algebra automorphism, and for an odd operator $\delta$ define its Laplacian $\Delta_{\delta}=\delta^{*}\delta+\delta^{*}\delta$ with harmonics ${\mathcal{H}}^{k}_{\delta}={\rm Ker}\,\Delta_{\delta}\cap\Omega^{k}$, where $\Omega^{k}$ is the space of real differential $k$-forms.
Lemma 2.
Let $(M,\omega)$ be a compact symplectic manifold of dimension $2n$. The following are equivalent:
  1. (1)
    The hard Lefschetz condition holds in degree $1$.
  2. (2)
    For any compatible almost-Kähler structure, ${\mathcal{H}}^{1}_{d}={\mathcal{H}}^{1}_{d^{c}}$, i.e. every $d$-harmonic $1$-form is $d^{c}$-harmonic (and vice versa).
  3. (3)
    For any compatible almost-Kähler structure, every $d$-harmonic $1$-form is the sum of $d$-harmonic forms of pure bi-degree,
    ${\mathcal{H}}^{1}_{d}={\mathcal{H}}^{1,0}_{d}\oplus{\mathcal{H}}^{0,1}_{d}.$
  4. (4)
    For any compatible almost-Kähler structure,
    ${\mathcal{H}}^{1}_{d}={\mathcal{H}}^{1}_{\bar{\partial}}\cap{\mathcal{H}}^{1}_{\mu},$
    where $d={\bar{\mu}}+{\bar{\partial}}+\partial+\mu$, with bidegrees
    $|{\bar{\mu}}|=(-1,2),\,|{\bar{\partial}}|=(0,1),\,|\partial|=(1,0),\text{ and }\,|\mu|=(2,-1).$
A proof that the first condition implies the second is given in  [14, Proposition 3.3], which we streamline below, using the same essential idea. We note that in  [19, Theorem 5.4], Tomassini and Wang show that the second condition implies first (in fact, in all degrees). Namely, for an almost-Kähler manifold, the space ${\mathcal{H}}_{d}$ satisfies the hard Lefschetz condition whenever ${\mathcal{H}}_{d}={\mathcal{H}}_{d^{c}}$ (in all degrees). The remaining statements are included to give additional characterizations in terms of decompositions of $d$-harmonic forms.
Proof.
To prove the first implication, assume there’s a $d$-harmonic 1-form $\alpha$ so that $J\alpha$ is not $d$-harmonic. Since $J$ is orthogonal, we may as well assume $\alpha$ is $d$-harmonic, but $J\alpha$ is orthogonal to the $d$-harmonics. Then
$L^{n-1}\alpha=\frac{1}{(n-1)!}\star J\alpha.$
Since star is an isometry, $L^{n-1}\alpha$ is orthogonal to the $d$-harmonics, and therefore $[L^{n-1}\alpha]$ is zero in cohomology. This shows $L^{n-1}$ has a non-zero kernel.
Using
$d^{c}$ $=-i{\bar{\mu}}+i{\bar{\partial}}-i\partial+i\mu$
we have
${\mathcal{H}}_{d}^{1}\cap{\mathcal{H}}_{d^{c}}^{1}={\mathcal{H}}^{1,0}_{d}\oplus{\mathcal{H}}^{0,1}_{d}\subset{\mathcal{H}}^{1}_{d}.$
This proves the second condition implies the third.
Next,
${\mathcal{H}}^{1,0}_{d}\oplus{\mathcal{H}}^{0,1}_{d}\subset{\mathcal{H}}^{1}_{\bar{\partial}}\cap{\mathcal{H}}^{1}_{\mu}\subset{\mathcal{H}}^{1}_{d}$
where the last containment follows from the almost-Kähler identity [8],
$\Delta_{\bar{\partial}}+\Delta_{\mu}=\Delta_{\partial}+\Delta_{\bar{\mu}}.$
This proves the third condition implies the fourth.
Finally, by  [8, Theorem 5.1], the last condition implies the first, since the space ¯${\mathcal{H}}_{\bar{\partial}}\cap{\mathcal{H}}_{\mu}$ always has the Lefschetz $\mathfrak{sl}(2)$-representation, and so satisfies hard Lefschetz on $H^{1}_{d}(M)\cong{\mathcal{H}}^{1}_{d}(M)$ if ¯${\mathcal{H}}^{1}_{\bar{\partial}}\cap{\mathcal{H}}^{1}_{\mu}={\mathcal{H}}^{1}_{d}(M)$. ∎
The proof of Theorem 1 is now essentially the same as in Vaisman [20], see also Ornea and Verbitsky [17].
Proof.
Let $(\eta,\theta)$ be a locally conformally symplectic structure. Suppose there is a symplectic structure $(M,\omega)$ satisfying hard Lefschetz in degree $1$, and let
$\theta=h+d\gamma$
be the Hodge decomposition of $\theta$, with $h\in{\mathcal{H}}_{d}^{1}$. By the conformal change $\eta\mapsto\eta^{\prime}=e^{-\gamma}\eta$, we have
$d\eta^{\prime}=\theta^{\prime}\wedge\eta^{\prime}$
and $\theta^{\prime}=h$ satisfies $d\theta^{\prime}=0$. So, we may as well assume that $d\eta=\theta\wedge\eta$ with $\theta$ $d$-harmonic.
By the Lemma’s first implication, $d^{c}\theta=0$ for any compatible almost Kähler structure, so that
$dd^{c}\eta^{n-1}=-(n-1)^{2}\theta\wedge J\theta\wedge\eta^{n-1},$
since $d^{c}\eta=J^{-1}d\eta=J^{-1}(\theta\wedge\eta)=-J\theta\wedge\eta$. Therefore,
[…]
Since $J$ is compatible with $\eta$, we conclude that $\theta=0$, so that $\eta$ is symplectic. ∎
Example 3.
A locally conformally symplectic structure $(\eta,\theta)$ on a symplectic manifold $(M,\omega)$ can indeed fail to be globally conformally symplectic if $\omega$ and $\eta$ do not share a compatible $J$. For example, the nilmanifold with structure equations $dx_{1}=x_{1}x_{3}$, $dx_{2}=-x_{2}x_{3}$, and $dx_{3}=dx_{4}=0$, with symplectic form $\omega=x_{1}x_{2}+x_{3}x_{4}$ has $(\theta=x_{3},\eta=x_{1}x_{4}+x_{2}x_{3})$ which is locally conformally symplectic, but not globally conformally symplectic. Note that $\omega$ does satisfy the hard Lefschetz in degree $1$.
On the other hand, the nilmanifold with structure equations $dx_{i}=x_{i-1}x_{1}$ for $i=3,4$, has locally conformally symplectic data ($\theta=x_{2},\eta=x_{1}x_{3}+x_{2}x_{4})$, which is not globally conformally symplectic, since $x_{2}\in H^{1}_{d}$ is non-trivial. Note that $\omega=x_{1}x_{4}+x_{2}x_{3}$ is symplectic, but does not satisfy hard Lefschetz in degree $1$. In this example, $\omega$ and $\eta$ do not share a compatible $J$.
We do not know of an example where $\omega$ is symplectic, $\eta$ is locally conformally symplectic but not globally conformally symplectic, while $\eta$ and $\omega$ share a compatible $J$. By Theorem 1, any such example could not satisfy the hard Lefschetz condition in degree $1$, and by Theorem 7 below, there are no examples on solvmanifolds.
From the theorem we deduce several corollaries, and emphasize that those statements concerning homogeneous spaces apply to all symplectic and locally conformally symplectic structures (even inhomogeneous structures).
Corollary 4.
Let $G$ be a simply connected $4$-dimensional Lie group admitting a uniform lattice $\Gamma$. Suppose $M=\Gamma\backslash G$ admits a symplectic form $\omega$ that satisfies hard Lefschetz in degree $1$. Let $(\eta,\theta)$ be a locally conformally symplectic structure on $M$. Suppose that there is an almost complex structure $J$ compatible with $\eta$ that is compatible with some symplectic form. Then, $(\eta,\theta)$ is a globally conformal symplectic structure.
Proof.
By hypothesis, $\omega$ satisfies hard Lefschetz in degree $1$, and therefore all degrees. By  [18, Theorem 4.2], any other symplectic form satisfies hard Lefschetz as well, so the corollary follows then from Theorem 1. ∎
Corollary 5.
Let $M$ be a $2$-dimensional complex torus or a hyperelliptic surface. Let $(\eta,\theta)$ be a locally conformally symplectic structure on $M$. Suppose that there is an almost complex structure $J$ compatible with $\eta$ that is compatible with some symplectic form. Then, $(\eta,\theta)$ is a globally conformal symplectic structure.
Proof.
It follows from [5, 6] that any symplectic form on a $2$-dimensional complex torus or a hyperelliptic surface satisfies hard Lefschetz. We also refer the reader to [13, Proposition 4.10] where it is proved that the cohomology Kähler cone coincides with the symplectic cone on a $2$-dimensional complex torus (see also [11]). The corollary then follows from Theorem 1. ∎
Tan and Tomassini studied in [18] the hard Lefschetz condition on some $6$-dimensional compact completely solvable manifolds namely Nakamura manifolds $N^{6}$ [16], and $M^{6}(c),N^{6}(c),P^{6}(c)$ (we refer the reader to [9, 12, 18] for their definition). We remark that the manifolds $M^{6}(c),N^{6}(c)$, and $P^{6}(c)$ do not admit any Kähler metric [12].
Corollary 6.
Let $M$ be a Nakamura manifold $N^{6}$ or the compact completely solvable manifold $M^{6}(c)$ or $N^{6}(c)$ or $P^{6}(c)$. Let $(\eta,\theta)$ be a locally conformally symplectic structure on $M$. Suppose there is an almost complex structure $J$ compatible with $\eta$ that is compatible with some symplectic form. Then, $(\eta,\theta)$ is a globally conformal symplectic structure.
Proof.
It follows from [18, Theorem 5.1, Theorem 6.1, Remark 6.2] that any symplectic form on $M$ satisfies hard Lefschetz. The corollary then follows from Theorem 1. ∎
The next theorem shows in particular that on compact solvmanifolds $\Gamma\backslash G$ with left-invariant structures, the hard Lefschetz assumption of Theorem 1 can be dropped. We recall that in such cases, the existence of a uniform lattice $\Gamma$ in a simply connected Lie group $G$ implies $G$ is unimodular [15], and if $\Gamma\backslash G$ is symplectic, then $G$ is solvable [7].
Theorem 7.
Let $\mathfrak{g}$ be a $2n$-dimensional unimodular Lie algebra with $\omega\in\bigwedge^{2}\mathfrak{g}^{*}$ satisfying $d\omega=0$. If $\eta\in\bigwedge^{2}\mathfrak{g}^{*}$, and $\omega$ and $\eta$ share a compatible $J$ on $\mathfrak{g}$, then $d\eta=0$.
In particular if $(\eta,\theta)$ is a (linear) locally conformally symplectic structure, and $\omega$ and $\eta$ share a compatible $J$, then $\eta$ is a symplectic.
Note the last example in Example 3 is unimodular, so the assumption that $\omega$ and $\eta$ share a compatible $J$ is essential.
Proof.
We denote by $h$ the almost-Kähler metric induced by $(\omega,J)$ and by $g$ the almost-Hermitian metric induced by $(\eta,J)$. Then, we can diagonalize $g$ with respect to the metric $h$ so there exists a $h$-orthonormal basis $\{e^{1},e^{2}=Je^{1},\ldots,e^{2n-1},e^{2n}=Je^{2n-1}\}$ of the dual of $\mathfrak{g}$ so that
$g=\sum_{i=1}^{n}\lambda_{i}\left(e^{2i-1}\otimes e^{2i-1}+e^{2i}\otimes e^{2i}\right),$
and so
$\eta=\sum_{i=1}^{n}\lambda_{i}\left(e^{2i-1}\wedge e^{2i}\right),$
where $\lambda_{i}$ are constant. On the other hand $\omega=\sum_{i=1}^{n}e^{2i-1}\wedge e^{2i}$. Hence,
$\sum_{i=1}^{n}d\left(e^{2i-1}\wedge e^{2i}\right)=0.$
The unimodular condition will be shown to imply that each summand above vanishes. Denote by $\{e_{i}\}$ the basis of $\mathfrak{g}$. We compute
$d\left(e^{2i-1}\wedge e^{2i}\right)(e_{2i-1},e_{2i},e_{2j-1}),\quad d\left(e^{2j-1}\wedge e^{2j}\right)(e_{2i-1},e_{2i},e_{2j-1}),$
$d\left(e^{2i-1}\wedge e^{2i}\right)(e_{2i-1},e_{2i},e_{2j}),\quad d\left(e^{2j-1}\wedge e^{2j}\right)(e_{2i-1},e_{2i},e_{2j}),$
for all indices $1\leq i,j\leq n$ with $i\neq j.$ For instance, for $i\neq j$,
(1) $d\left(e^{2i-1}\wedge e^{2i}\right)(e_{2i-1},e_{2i},e_{2j-1})=e^{2i-1}\left([e_{2i-1},e_{2j-1}]\right)+e^{2i}\left([e_{2i},e_{2j-1}]\right),$
and
(2) $d\left(e^{2j-1}\wedge e^{2j}\right)(e_{2i-1},e_{2i},e_{2j-1})=e^{2j}\left([e_{2i-1},e_{2i}]\right).$
Since $d\omega=0,$ we get from Equations (1) and (2) that
$e^{2i-1}\left([e_{2i-1},e_{2j-1}]\right)+e^{2i}\left([e_{2i},e_{2j-1}]\right)+e^{2j}\left([e_{2i-1},e_{2i}]\right)=0.$
Fixing $j$ and summing over $1\leq i\leq n$ with $i\neq j$, we get
$\sum_{i\neq j}e^{2i-1}\left([e_{2i-1},e_{2j-1}]\right)+\sum_{i\neq j}e^{2i}\left([e_{2i},e_{2j-1}]\right)+\sum_{i\neq j}e^{2j}\left([e_{2i-1},e_{2i}]\right)=0.$
Using the fact that $\mathfrak{g}$ is unimodular,
$e^{2j}\left([e_{2j-1},e_{2j}]\right)+\sum_{i\neq j}e^{2j}\left([e_{2i-1},e_{2i}]\right)=0,$
so that
$e^{2j}\left(\sum_{i=1}^{n}[e_{2i-1},e_{2i}]\right)=\sum_{i=1}^{n}e^{2j}\left([e_{2i-1},e_{2i}]\right)=0,$
for all $1\leq j\leq n$. Equivalently, for all $1\leq j\leq n$
$e^{2j-1}\left(\sum_{i=1}^{n}[e_{2i-1},e_{2i}]\right)=0.$
It follows that for each $i$, $\operatorname{span}\{e_{2i-1},e_{2i}\}$ is a Lie subalgebra of $\mathfrak{g}$ and therefore $d\left(e^{2i-1}\wedge e^{2i}\right)=0$ for all $i=1,\ldots,n$. Therefore, $\eta$ is a symplectic form. ∎
In the non-invariant case, the follow question alluded to in Example 3 seems to be open:
Question 8.
Is there a compact symplectic manifold $(M,\omega)$ with a locally conformally symplectic structure $(\eta,\theta)$, which is not globally conformally symplectic, such that $\eta$ and $\omega$ share a compatible $J$?
Acknowledgements
The authors are grateful to Giovanni Bazzoni and Liviu Ornea for useful discussions about Question 8.